Determine whether the relation R on the set of all real

numbers is reflexive, symmetric, antisymmetric, and/or

transitive, where (x, y) ∈ R if and only if

a) x + y = 0.

b) x = ±y.

c) x - y is a rational number.


1
Expert's answer
2021-09-27T12:28:28-0400

Part a

Consider the relation "R : x+ y = 0, for \space x,y \in R" .

Reflexive:

For "1 \in R, 1+1=2 \not=0."

Therefore, R is not reflexive.

Symmetric:

For "x,y \in R, x+y=0=y+x."

Therefore, R is symmetric.

Anti-symmetric:

For "-1,1 \in R, -1+1=0 \space and \space 1+(-1)=0 \space but \space -1 \not=1"

Therefore, R is not anti-symmetric.

Transitive:

"For -1,1 \in R, 1+(-1)=0 \space and \space (-1)+1=0"

That is "(1,-1) \in R" and "(-1,1) \in R" but "(1,1) \not \in R"

Therefore, R is not transitive. 


Part b

Consider the relation "R : ±y = 0, for \space x,y \in R" .

Reflexive:

For "x \in R, x=x \implies x=±x, so (x,x) \in R"

Therefore, R is reflexive.

Symmetric:

For "x,y \in R, x=±y \implies y=±x."

Therefore, R is symmetric.

Anti-symmetric:

For "(-1,1) \in R, (1,-1) \in R; -1 = ±1 \space and \space 1=±(-1) \space but \space -1 \not=1"

Therefore, R is not anti-symmetric.

Transitive:

"For (x,y) \in R, (y,z) \in R \\
x= ±y \space and \space y=±z"

That is "x= \u00b1z"

Therefore, R is transitive. 


Part c

Consider the relation "R:x-y" is a rational number for "x,y \in R"

Reflexive: For "x \in R, x-x=0" is a rational number, so "(x,x) \in R" .

Therefore, R is reflexive.

Symmetric: For "(x,y) \in R" ,

"x-y" is rational

"y -x" is rational

"(y,x) \in R"

Therefore, R is symmetric.

Anti-symmetric:

For "(1,2) \in R" and "(2,1) \in R" ,

That is, 1— 2 is rational and 2 is rational, but "2 \not =1" .

Therefore, R is not anti-symmetric.

Transitive:

For "(x,y) \in R" and "(y,z) \in R", then x— y is rational and y—z is rational

Implies that x—z=x—y+y—z

As the sum of two rational numbers is rational follows that x — z is rational.

Therefore, R is transitive.